Two immiscible liquids of relative densities 1.5 and 2 are at rest in a rectangular container of dimensions 3×2×1m.Find the forCe exerted on the bottom of the container

Grade:11

1 Answers

deepak
7 years ago
the deinsity of the first liquid = 1.5 g/cm3
the density of the second liquid = 2 g/cm3
length of the container = 300cm
breath of the container = 200cm
depth of the container =100 cm
let the first liquid befilled to a hieght of h(cm)
the the hieght of liquid 2 = 100−h
then, mass of liquid one stored in the container = 600hx1.5=900g=0.9Kg
similarly mass of liq two stored = 600(100−h)x2=120000−1200h g=120−1.2h Kg
therefore force due to gravity on the two mases at the base=(120−0.3h)9.8 N
 
for finding the force in any configuration just substitute the value of h in cm
hope this proves to be helpful.  XP